a) How much current does your laptop consume if 100 C of charge is transferred through the charger wire in 4 seconds? 

b) How much time is required for a 20 Coulombs charge to flow through a point if the current is 2 amperes? 



1
Expert's answer
2021-11-03T11:12:24-0400

"Part\>a) \\\\\\
\textcolor{blue}{\Iota=\frac{Q}{t}}" Where:


"\Iota" = current

"Q" = Charge

"t" = time


"\therefore\>\Iota=\frac{100}{4}=25\>A"


Part b)


"\textcolor{blue}{t=\frac{Q}{\Iota}}"


"t=\frac{20}{2}=10\>second"